# MRGen: Segmentation Data Engine For Underrepresented MRI Modalities

### v2

@@ Abstract
Training medical image segmentation models for rare yet clinically significant imaging modalities is challenging due to the scarcity of annotated data, and manual mask annotations can be costly and labor-intensive to acquire. This paper investigates leveraging generative models to synthesize training data, to train segmentation models for underrepresented modalities, particularly on annotation-scarce MRI. Concretely, our contributions are threefold: (i) we introduce MRGen-DB, a large-scale radiology image-text dataset comprising extensive samples with rich metadata, including modality labels, attributes, regions, and organs information, with a subset having pixelwise mask annotations; (ii) we present MRGen, a diffusion-based data engine for controllable medical image synthesis, conditioned on text prompts and segmentation masks. MRGen can generate realistic images for diverse MRI modalities lacking mask annotations, facilitating segmentation training in low-source domains; (iii) extensive experiments across multiple modalities demonstrate that MRGen significantly improves segmentation performance on unannotated modalities by providing high-quality synthetic data. We believe that our method bridges a critical gap in medical image analysis, extending segmentation capabilities to scenarios that are challenging to acquire manual annotations. The codes, models, and data will be publicly available at https://haoningwu3639.github.io/MRGen/.
@@ 1 Introduction
In this paper, we investigate the potential of generative models, particularly diffusion models, to synthesize MRI data for training segmentation models on underrepresented modalities. While generative models offer a promising solution, they face unique challenges: (i) data availability remains a significant obstacle. Existing approaches have primarily focused on data augmentation for well-annotated modalities such as X-ray [3] and CT [14, 15], as depicted in Figure 1 (right). However, MRI data is relatively scarce, and highly diverse across modalities, making it less explored for generative modeling; (ii) controllability is critical to facilitate downstream tasks such as segmentation. Thus, generative models must enable controllable synthesis based on conditions such as texts, masks, or both. Yet, prior works [55, 9, 42, 60, 32] cannot simultaneously support both conditions, limiting their abilitiy to control the generated modalities, regions, and organs effectively.
@@
Our first contribution is the collection of a large-scale, high-quality radiology image-text dataset, MRGen-DB (short for “Database for MRI Generation”), which includes MRI scans across various modalities sourced from the Internet and open-source repositories. The dataset consists of nearly 250,000 2D slices, enriched with detailed annotations such as modality labels, attributes, region, and organ information, with a subset providing organ masks. This extensive collection of image-text pairs across diverse modalities forms a robust foundation for training a general MRI generation model, while the availability of mask annotations facilitates more controllable and targeted synthesis.
@@
For controllable data generation, we present MRGen, a diffusion-based data engine for MRI synthesis, that supports conditioning on both text prompts and segmentation masks. We employ a two-stage training strategy: (i) text-guided pretraining on diverse, large-scale image-text pairs, enabling the model to synthesize images across various modalities guided by templated text descriptions; and (ii) mask-conditioned finetuning on mask-annotated data, facilitating controllable generation based on organ masks. Consequently, such a two-stage strategy allows MRGen to extend its controllable generation abilities towards modalities that originally do not have segmentation annotations available, thereby enables training segmentation models for these underrepresented modalities with synthetic data.
@@
Overall, our contributions can be summarized as follows: (i) we explore the use of generative models for MRI synthesis across annotation-scarce modalities, facilitating training segmentation models for underrepresented modalities; (ii) we curate MRGen-DB, a large-scale radiology image-text dataset, which features detailed modality labels, attributes, regions, and organs information, with a subset of organ mask annotations, providing a robust foundation for medical generative modeling; (iii) we develop MRGen, a diffusion-based data engine capable of controllable generation, conditioned on templated text prompts and segmentation masks; (iv) we conduct extensive experiments across diverse modalities, demonstrating that MRGen can controllably generate high-quality MR images, improving ‘zero-shot’ segmentation performance for unannotated modalities. To our knowledge, this work introduces the first open-source dataset curated for medical image generation, and the first general medical generative model tailored for annotation-scarce MRI modalities, offering a novel solution to address the scarcity of medical data and annotations.
2 Related works
@@
Figure 2: Architecture Overview. Developing our MRGen involves three key steps: (a) Train an autoencoder on various images from dataset $\mathcal{D}_{u}$; (b) Train a text-guided generative model within the latent space, using image-text pairs across diverse modalities from MRGen-DB, featuring modality, attributes, region, and organs information; (c) Train a mask condition controller jointly on image-text pairs with and without mask annotations, enabling controllable generation based on both text prompts and masks.
@@ 3 Method
Here, we start to formulate the problem of interest in Sec. 3.1, followed by a detailed description of the dataset curation procedure in Sec. 3.2. Later, we elaborate on the proposed MRGen architecture and the training details of our model in Sec. 3.3 and Sec. 3.4, respectively. Lastly, we present the procedure of synthesizing and filtering samples for downstream segmentation tasks in Sec. 3.5.
@@ 3.1 Problem Formulation
where $\Theta$ and $\Theta_{c}$ refer to parameters of the generative model and the mask condition controller, respectively. Developing such a controllable data engine, thus enables synthesizing high-quality data to train segmentation models for the challenging ‘zero-shot’ scenario.
Relations to existing tasks. Numerous studies have proven the effectiveness of generative models for data augmentation [55] on well-annotated modalities and regions, such as CT [15, 14], X-ray [3], and brain MRI [9, 42], however, this is not the focus of our work. Instead, we target the more challenging scenario of synthesizing MR images for scarce and underrepresented modalities where no manual mask annotations are available. While image translation methods [69, 27, 52, 46] offer a potential alternative by translating richly annotated data to underrepresented modalities, these approaches often require registered data for training or are limited to specific modality conversions. In contrast, our proposed MRGen framework offers a flexible and controllable generation pipeline, enabling the synthesis of complex abdominal MRI data across diverse modalities, even in the absence of mask annotations.
@@ 3.2 Dataset Curation
Automatic annotations. Abdominal imaging is highly variable, exhibiting significant differences across anatomical regions, such as the Upper Abdominal Region and the Pelvic Region. Relying solely on modality labels is insufficient to differentiate these distinctions. To address this, we divide the abdomen into six anatomical regions: Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region. Using the pre-trained BiomedCLIP model [62], we automatically categorize all 2D slices into these regions. To maintain annotation quality, slices with low confidence scores ($<40\%$) are intentionally left unlabeled. This process enriches the dataset with detailed region-specific information.
@@
To ensure the reliability of the automatic annotations, we have uniformly sampled and manually verified a subset of the data. Specifically, 2% of the region annotations and 20% of the modality attributes have been reviewed, achieving high accuracies of 95.33% and 91.67%, respectively. This verification step ensures the high quality of the dataset and strengthens its applicability for downstream tasks.
Discussion. After completing the data processing above, we assemble the MRGen-DB dataset, which includes approximately 6,000 3D volumes spanning over 100 distinct free-text MR modalities, totaling nearly 250,000 2D slices, as presented in Table 1. Each sample is paired with its modality label, attributes, region, and organ information, with about 18,000 samples featuring mask annotations. The scale, diversity, and fine-grained annotations of MRGen-DB provide sufficient information for training generative models tailored for MR images. More statistics are provided in Sec. B.2.
@@ 3.3 Architecture
To enable the generative model to effectively learn controllable generation based on texts and masks, the training process is carried out in two stages: (i) pretraining a text-guided generative model on image-text data, covering diverse modalities; and (ii) finetuning a mask condition controller jointly on data with and without mask annotations.
@@
As depicted in Figure 2 (b), the denoising process adopts a UNet [50] and reconstructs images from noise by estimating the noise term $\hat{\boldsymbol{\epsilon}}$. Concretely, to generate images guided by text prompts, we design templated text prompt ($\mathcal{T}$), that consists of diverse modality labels, modality attributes, regions, and organs information, for example: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] “T1 MRI; fat high signal, muscle intermediate signal, water low signal, fat bright, water dark; upper abdomen; liver, spleen, and kidney”. These templated prompts ensure sufficient clinical information to distinguish distinct modalities, regions, and organs. We employ an off-the-shelf BiomedCLIP [62] text encoder ($\phi_{\mathrm{text}}$) to encode them into embeddings, denoted as $\mathcal{C}_{T}$ = $\phi_{\mathrm{text}}(\mathcal{T})$. These embeddings are integrated into our model via cross-attention, serving as the key and value, with visual features ($\mathbf{z}_{t}$) as the query. The output ($\mathbf{O}_{\mathrm{cross}}$) of each cross-attention layer ($\mathcal{F}_{\mathrm{cross}}$) is represented as:
@@ 3.4 Model Training
Autoencoder reconstruction. The autoencoder for compression is trained on raw images from MRGen-DB, using a combination of MSE loss and KL divergence loss as follows: $\mathcal{L}_{\mathrm{VAE}}=||\mathcal{I}-\hat{\mathcal{I}}||_{2}^{2}+\gamma\mathcal{L}_{\mathrm{KL}}$, where $\mathcal{L}_{KL}$ imposes a KL-penalty towards a standard normal on the learned latent, similar to VAE [29] and $\gamma$ denotes a predefined weight.
@@
Figure 3: Synthetic Data Construction Pipeline. MRGen takes text prompt and mask as conditions for controllably generating MR images and employs a pretrained SAM2 model for automatic filtering to guarantee the quality of generated samples.
@@ 3.5 Synthetic Data for Segmentation Training
With our data engine, we can then produce MR samples for training downstream segmentation models. At inference time, the text prompt ($\mathcal{T^{\prime}}$) and controlling organ mask ($\mathcal{M^{\prime}}$) are fed into our MRGen model ($\Phi_{\mathrm{MRGen}}$) as conditions to generate the corresponding MR sample ($\mathcal{I^{\prime}}$). To ensure the fidelity of the generated images on mask conditions, we design an automatic filtering pipeline using the off-the-shelf SAM2-Large [48] model, as depicted in Figure 3. Specifically, we feed the conditional mask ($\mathcal{M^{\prime}}$) and the generated image ($\mathcal{I^{\prime}}$) into SAM2 to predict a segmentation map with a confidence score ($s_{\mathrm{conf}}$), which is used to calculate the IoU score ($s_{\mathrm{IoU}}$) against $\mathcal{M^{\prime}}$. A sample is considered to be high-quality and aligned with the mask condition if both its IoU score ($s_{\mathrm{IoU}}$) and confidence score ($s_{\mathrm{conf}}$) exceed the predefined thresholds; otherwise, it is discarded.
@@ 4.1 Experimental Settings
Unlike existing work that focuses on data augmentation for well-annotated modalities, we explore the more challenging scenarios where certain modalities lack annotations entirely, and aim to leverage generative models to synthesize data for training segmentation models towards these underrepresented modalities. Specifically, to simulate such a clinical scenario, we construct 5 cross-modality dataset pairs within our MRGen-DB, each comprising a mask-annotated source-domain dataset ($\mathcal{D}_{s}$) and an unannotated target-domain dataset ($\mathcal{D}_{t}$). Models trained on each dataset pair synthesize target-domain samples for training segmentation models. We assess our data engine from two aspects: (i) image generation quality and (ii) downstream segmentation performance on the target-domain test set.
@@
Baselines. For generation, we compare generated images from MRGen against three representative approaches: CycleGAN [69] and UNSB [27] for translating source-domain images to the target domain; and DualNorm [68] for exhaustive augmentation of source-domain images. For segmentation, we evaluate models trained on five data sources: (i) source-domain data, as a baseline; (ii) DualNorm augmented data; (iii) CycleGAN translated data; (iv) UNSB translated data; and (v) MRGen generated data. We adopt nnUNet [23] and UMamba [40] as segmentation frameworks for all settings, except for DualNorm, which employs a customized UNet following their official codes. Comparisons with additional baselines are provided in Sec. D.2.
Evaluation metrics. We employ distinct metrics for the evaluation: For image generation, we employ Fréchet Inception Distance (FID) [18] score to assess the diversity and quality of generated images. For segmentation models, we employ commonly used Dice Similarity Coefficient (DSC) [41] score to compare predicted masks with ground truth. Considering segmentation consistency, we stack slices into 3D volumes, calculate the DSC for each organ individually, and average them as the final result.
Implementation details. All images are resized to 512 $\times$ 512, and training is conducted on 8×8$\times$8 × Nvidia A100 GPUs using the AdamW [38] optimizer. We start by training the autoencoder with a learning rate of $5\times 10^{-5}$ and a batch size of $256$ for 50K iterations. Next, the text-guided generative model and mask condition controller are trained with a learning rate of $1\times 10^{-5}$, using batch sizes of 256 and 128 for 200K and 40K iterations, respectively. Moreover, we randomly drop text prompts with a 10% probability to enable classifier-free guidance [19]. The compression ratio, latent dimension $d$, KL loss weight $\gamma$, and diffusion timesteps $T$ are set to $8$, $16$, $1\times 10^{-4}$, and $1000$, respectively. During inference, we perform 50-step sampling using DDIM [53], with classifier-free guidance weight $w=7.0$. For each mask, we generate 20 image candidates and select the best two satisfying the predefined thresholds, which are set to 0.80 and 0.90 for IoU and confidence scores, respectively. Conditions for target-domain data synthesis are directly derived from region, organs information, and segmentation masks of source-domain data.

### v3

@@ Abstract
Training medical image segmentation models for rare yet clinically important imaging modalities is challenging due to the scarcity of annotated data, and manual mask annotations can be costly and labor-intensive to acquire. This paper investigates leveraging generative models to synthesize data, for training segmentation models for underrepresented modalities, particularly on annotation-scarce MRI. Concretely, our contributions are threefold: (i) we introduce MRGen-DB, a large-scale radiology image-text dataset comprising extensive samples with rich metadata, including modality labels, attributes, regions, and organs information, with a subset featuring pixel-wise mask annotations; (ii) we present MRGen, a diffusion-based data engine for controllable medical image synthesis, conditioned on text prompts and segmentation masks. MRGen can generate realistic images for diverse MRI modalities lacking mask annotations, facilitating segmentation training in low-source domains; (iii) extensive experiments across multiple modalities demonstrate that MRGen significantly improves segmentation performance on unannotated modalities by providing high-quality synthetic data. We believe that our method bridges a critical gap in medical image analysis, extending segmentation capabilities to scenarios that are challenging to acquire manual annotations. The codes, models, and data will be publicly available at https://haoningwu3639.github.io/MRGen/.
@@ 1 Introduction
In this paper, we investigate the potential of generative models, particularly diffusion models, to synthesize MRI data for training segmentation models on underrepresented modalities. While generative models offer a promising solution, they face unique challenges: (i) data availability remains a significant obstacle. Existing approaches have primarily focused on data augmentation for well-annotated modalities such as X-ray [4] and CT [14, 15], as depicted in Figure 1 (right). However, MRI data is relatively scarce, and highly diverse across modalities, making it less explored for generative modeling; (ii) controllability is critical to facilitate downstream tasks such as segmentation. Thus, generative models must enable controllable synthesis based on conditions such as texts, masks, or both. Yet, prior works [56, 9, 43, 61, 33] cannot simultaneously support both conditions, limiting their ability to control the generated modalities, regions, and organs effectively.
@@
Our first contribution is to curate a high-quality, large-scale radiology image-text dataset, termed as MRGen-DB, which includes MRI scans across various modalities sourced from the Internet and open-source repositories. The dataset consists of nearly 250,000 2D slices, enriched with detailed annotations such as modality labels, attributes, region, and organ information, with a subset providing organ masks. This extensive collection of image-text pairs across diverse modalities forms a robust foundation for training an MRI generative model, while the availability of mask annotations facilitates more controllable and targeted synthesis.
@@
For controllable data generation, we present MRGen, a diffusion-based data engine for MRI synthesis, that supports conditioning on both text prompts and segmentation masks. We employ a two-stage training strategy: (i) text-guided pretraining on diverse, large-scale image-text pairs, enabling the model to synthesize images across various modalities guided by templated text descriptions; and (ii) conditioned finetuning on mask-annotated data, facilitating controllable generation based on organ masks. Consequently, such a two-stage strategy allows MRGen to extend its controllable generation abilities towards modalities that originally do not have segmentation annotations available, thereby enabling training segmentation models for the underrepresented modalities with synthetic data.
@@
Overall, our contributions can be summarized as follows: (i) we explore the use of generative models for MRI synthesis across annotation-scarce modalities, facilitating training segmentation models for underrepresented modalities; (ii) we curate MRGen-DB, a large-scale radiology image-text dataset, which features detailed modality labels, attributes, regions, and organs information, with a subset of organ mask annotations, providing a robust foundation for medical generative modeling; (iii) we develop MRGen, a diffusion-based data engine capable of controllable generation, conditioned on templated text prompts and segmentation masks; (iv) we conduct extensive experiments across diverse modalities, demonstrating that MRGen can controllably generate high-quality MR images, improving ‘zero-shot’ segmentation performance for unannotated modalities. To the best of our knowledge, this work introduces the first open-source dataset curated for medical image generation, and the first generative model tailored for general annotation-scarce MRI modalities, offering a novel solution to address the scarcity of medical data and annotations.
2 Related Work
@@
Figure 2: Architecture Overview. Developing our MRGen involves three key steps: (a) Train an autoencoder on various images from MRGen-DB dataset; (b) Train a text-guided generative model within the latent space, using image-text pairs across diverse modalities from MRGen-DB, featuring modality, attributes, region, and organs information; (c) Train a mask condition controller jointly on image-text pairs with and without mask annotations, enabling controllable generation based on both text prompts and masks.
@@ 3 Method
We start to formulate the problem of interest in Sec. 3.1, followed by a detailed description of the dataset curation procedure in Sec. 3.2. Later, we elaborate on the proposed MRGen architecture and the training details of our model in Sec. 3.3 and Sec. 3.4, respectively. Lastly, we present the procedure of synthesizing and filtering samples for downstream segmentation tasks in Sec. 3.5.
@@ 3.1 Problem Formulation
where $\Theta$ and $\Theta_{c}$ refer to parameters of the generative model and the mask condition controller, respectively. Developing such a controllable data engine enables synthesizing high-quality data to train segmentation models for the challenging scenario that lacks manual annotations, for example, ‘zero-shot’ evaluation.
Relations to existing tasks. Numerous studies have proven the effectiveness of generative models for data augmentation [56] on well-annotated modalities and regions, such as CT [15, 14], X-ray [4], and brain MRI [9, 43], however, this is not the focus of our work. Instead, we target the more challenging scenario of synthesizing MR images for scarce and underrepresented modalities where no manual mask annotations are available. While image translation methods [70, 28, 53, 47] offer a potential alternative by translating richly annotated data to underrepresented modalities, these approaches often require registered data pairs for training or are limited to specific modality conversions. In contrast, our proposed MRGen framework offers a flexible and controllable generation pipeline, enabling the synthesis of complex abdominal MRI data across diverse modalities, even in the absence of mask annotations.
@@ 3.2 Dataset Curation
Automatic annotations. Abdominal imaging normally exhibits significant differences across anatomical regions, such as the Upper Abdominal Region and the Pelvic Region. Relying solely on modality labels is insufficient to differentiate these distinctions. To address this, we divide the abdomen into six anatomical regions: Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region. Using the pre-trained BiomedCLIP [63], we automatically classify all 2D slices into these regions. To maintain annotation quality, slices with low confidence ($<40\%$) are intentionally left unlabeled. This process enriches the dataset with detailed region-specific information.
@@
To ensure the reliability of the automatic annotations, we have uniformly sampled and manually verified a data subset. Specifically, 2% of the region annotations and 20% of the modality attributes have been reviewed, achieving high accuracies of 95.33% and 91.67%, respectively. This verification step ensures the high quality of the dataset and strengthens its applicability for downstream tasks.
Discussion. After the data processing above, we assemble the MRGen-DB dataset, which includes about 6,000 3D volumes spanning over 100 distinct free-text MR modalities, totaling nearly 250,000 2D slices, as shown in Table 1. In the following parts, we treat each 2D slice as a data sample, which is paired with its modality label, attributes, region, and organ information, with roughly 18,000 samples featuring mask annotations. The scale, diversity, and fine-grained annotations of MRGen-DB offer sufficient information for training generative models tailored for MR images. More statistics are provided in Sec. B.2 of the Appendix.
@@ 3.3 Architecture
To effectively learn controllable generation based on texts and masks, the training process is carried out in two stages: (i) pretraining a text-guided generative model on image-text data, covering diverse modalities; and (ii) finetuning a mask condition controller jointly on data with and without mask annotations.
@@
As depicted in Figure 2 (b), the denoising process adopts a UNet [51] and reconstructs images from noise by estimating the noise term $\hat{\boldsymbol{\epsilon}}$. Concretely, to generate images guided by text prompts, we design a templated text prompt ($\mathcal{T}$), that consists of diverse modality labels, modality attributes, regions, and organs information, for example: {mdframed}[backgroundcolor=gray!10, linewidth=0pt] “T1 MRI; fat high signal, muscle intermediate signal, water low signal, fat bright, water dark; upper abdomen; liver, spleen, and kidney”. These templated prompts ensure sufficient clinical information to distinguish distinct modalities, regions, and organs. We employ an off-the-shelf BiomedCLIP [63] text encoder ($\phi_{\mathrm{text}}$) to encode them into embeddings, denoted as $\mathcal{C}_{T}$ = $\phi_{\mathrm{text}}(\mathcal{T})$. These embeddings are integrated into our model via cross-attention, serving as the key and value, with visual features ($\mathbf{z}_{t}$) as the query. The output ($\mathbf{O}_{\mathrm{cross}}$) of each cross-attention layer ($\mathcal{F}_{\mathrm{cross}}$) is represented as:
@@ 3.4 Model Training
Autoencoder reconstruction. The autoencoder for compression is trained on raw images from MRGen-DB, using a combination of MSE loss and KL divergence loss as follows: $\mathcal{L}_{\mathrm{VAE}}=||\mathcal{I}-\hat{\mathcal{I}}||_{2}^{2}+\gamma\mathcal{L}_{\mathrm{KL}}$, where $\mathcal{L}_{KL}$ imposes a KL-penalty towards a standard normal on the learned latent, similar to VAE [30], and $\gamma$ denotes a predefined weight.
@@
Figure 3: Synthetic Data Construction Pipeline. MRGen takes a text prompt and a mask as conditions for controllably generating MR images and employs a pretrained SAM2 model for automatic filtering to guarantee the quality of generated samples.
@@ 3.5 Synthetic Data for Segmentation Training
With our data engine, we can then produce MR samples for training downstream segmentation models. At inference time, the text prompt ($\mathcal{T^{\prime}}$) and controlling organ mask ($\mathcal{M^{\prime}}$) are fed into our generative model ($\Phi_{\mathrm{MRGen}}$) as conditions to generate the corresponding MR sample ($\mathcal{I^{\prime}}$). To ensure the fidelity of the generated images on mask conditions, we design an automatic filtering pipeline using the off-the-shelf SAM2-Large [49] model, as depicted in Figure 3. Specifically, we feed the conditional mask ($\mathcal{M^{\prime}}$) and the generated image ($\mathcal{I^{\prime}}$) into SAM2 to predict a segmentation map with a confidence score ($s_{\mathrm{conf}}$), which is used to calculate the IoU score ($s_{\mathrm{IoU}}$) against $\mathcal{M^{\prime}}$. A sample is considered to be high-quality and aligned with the mask condition if both its IoU score ($s_{\mathrm{IoU}}$) and confidence score ($s_{\mathrm{conf}}$) exceed the predefined thresholds; otherwise, it is discarded.
@@ 4.1 Experimental Settings
Unlike existing work focusing on data augmentation for well-annotated modalities, we explore the more challenging scenario where certain modalities lack annotations entirely, and aim to adopt generative models to synthesize data for training segmentation models for such underrepresented modalities. Concretely, to simulate such a clinical scenario, we construct 5 cross-modality dataset pairs within MRGen-DB, each pair comprises a mask-annotated source-domain dataset ($\mathcal{D}_{s}$) and an unannotated target-domain dataset ($\mathcal{D}_{t}$). We leverage models trained on each dataset pair to generate target-domain samples for training segmentation models and assess our data engine from two aspects: (i) image generation quality and (ii) segmentation performance on the target-domain test set. Here, the test data for segmentation is strictly not used for training our generative data engines to prevent information leakage.
@@
Baselines. For generation, we compare generated images from MRGen against three representative methods: CycleGAN [70] and UNSB [28] for translating source-domain images to the target domain; and DualNorm [69] for exhaustive augmentation of source-domain images. For segmentation, we assess models trained on five data sources: (i) source-domain data, as a baseline; (ii) DualNorm augmented data; (iii) CycleGAN translated data; (iv) UNSB translated data; and (v) MRGen generated data. We adopt nnUNet [23] and UMamba [41] as segmentation frameworks for all settings, except for DualNorm, which employs a customized UNet following their official codes. More baseline results are provided in Sec. D.2 of the Appendix.
Evaluation metrics. For image generation, we leverage Fréchet Inception Distance (FID) [18] score to assess the diversity and quality of generated images. For segmentation models, we employ the commonly used Dice Similarity Coefficient (DSC) [42] score to compare predicted masks with ground truth. Considering segmentation consistency, we stack slices into 3D volumes, calculate the DSC for each organ individually, and average them as the final result.
Implementation details. All images are resized to 512 $\times$ 512, and training is conducted on 8×8$\times$8 × Nvidia A100 GPUs using the AdamW [39] optimizer. We start by training the autoencoder with a learning rate of $5\times 10^{-5}$ and a batch size of $256$ for 50K iterations. Next, the text-guided generative model and mask condition controller are trained with a learning rate of $1\times 10^{-5}$, using batch sizes of 256 and 128 for 200K and 40K iterations, respectively. Moreover, we randomly drop text prompts with a 10% probability to enable classifier-free guidance [19]. The compression ratio, latent dimension $d$, KL loss weight $\gamma$, and diffusion timesteps $T$ are set to $8$, $16$, $1\times 10^{-4}$, and $1000$, respectively. During inference, we perform 50-step sampling using DDIM [54], with classifier-free guidance weight $w=7.0$. For each mask, we generate 20 image candidates and select the best two satisfying the predefined thresholds, which are set to 0.80 and 0.90 for IoU and confidence scores, respectively. Conditions for target-domain data synthesis are directly derived from the region, organs information, and segmentation masks of the source-domain data.